# Deep Learning, Machine Learning - Digital Signal and Image Processing: From Theory to Application

### v1

@@ 7.1.2 Image Elements
The building blocks of any digital image are its pixels, which form a grid-like structure representing the image. Each pixel is a tiny square (or sometimes other shapes) that contains a value corresponding to the brightness (in grayscale images) or color (in color images) at that point in the image. The quality and detail of an image depend on two important factors: spatial resolution and grayscale resolution [58].
@@
  • In an 8-bit grayscale image, each pixel can have one of 256 possible intensity values (ranging from 0 to 255) [138].
@@
  • In a 24-bit color image, each pixel has 8 bits per channel (red, green, and blue), allowing for $216256^{3}=16,777,$ 216 different colors.
The higher the bit depth, the more shades or colors an image can represent, resulting in smoother gradients and more realistic visual representations [69].
@@ 7.2 Python Example: Working with Grayscale Images
Python offers excellent libraries for digital image processing, such as Pillow [36], Matplotlib [78], and OpenCV [21]. To begin with, let’s work with grayscale images, which are simpler to understand since they only contain intensity information.
@@ 7.4.1 Image Sampling
Image sampling is the process of converting a continuous image into a digital format by taking measurements, or samples, of the image at regular intervals. These samples correspond to pixels in the digital image, and each pixel represents the intensity (or color) of the image at a specific point [58].
In the continuous world, images are represented by a function $f(x,y)$, where $x$ and $y$ denote the spatial coordinates, and $f(x,y)$ represents the intensity at that point. When we sample the image, we select a grid of points at intervals and store the intensity values at those locations. This sampled image becomes a discrete grid of pixels, and the spacing between the samples is referred to as the sampling rate or resolution [179].
@@
The sampling rate, or the distance between adjacent samples, has a significant impact on the quality and detail of the digital image. A higher sampling rate (denser grid of pixels) captures more detail from the original image, while a lower sampling rate results in a loss of detail and potential aliasing effects [25].
  • High Sampling Rate: When the sampling rate is high, the image closely resembles the original continuous image. Fine details and sharp edges are preserved [58].
@@
The Nyquist Sampling Theorem provides a guideline for selecting an appropriate sampling rate to avoid aliasing [179]. According to this theorem, the sampling rate must be at least twice the highest frequency present in the image to accurately capture all details [136]. This minimum rate is known as the Nyquist rate. If the sampling rate is lower than the Nyquist rate, aliasing occurs, where high-frequency components appear as lower-frequency artifacts in the image [25].
@@ 7.4.2 Image Quantization
After sampling an image in the spatial domain, the next step is to quantize its intensity values. Quantization is the process of mapping the continuous range of intensity values in the original image to a limited set of discrete values in the digital image [58]. In practice, each pixel’s intensity in the digital image is represented by a fixed number of bits, which limits the number of possible intensity levels.
@@
  • 8-bit Image: Each pixel can have one of 256 possible intensity levels (from 0 to 255). This is common for grayscale images [86].
  • 16-bit Image: Each pixel can have one of 65,536 possible intensity levels (from 0 to 65535), allowing for much finer intensity gradations [62].
@@
Quantization introduces an error known as quantization error, which is the difference between the actual continuous intensity values and the discrete levels assigned during quantization. This error can manifest as noise or distortion in the image, especially when using a low bit depth [69].
@@ 7.5 Color Spaces
In digital image processing, color spaces are used to represent colors in a format that can be processed and analyzed by computers. Different color spaces are suited for different applications, from basic image display to more complex tasks like image compression, segmentation, and enhancement [58]. The choice of color space affects how easily certain operations can be performed on the image, and how colors are interpreted. In this section, we will cover several commonly used color models, including RGB, HSV/HSI, and YCbCr.
@@ 7.5.1 RGB Color Model
The RGB color model is one of the most widely used color models in digital image processing. It represents colors by combining three primary color channels: Red (R), Green (G), and Blue (B). Each pixel in an image has a value for these three channels, and by varying the intensity of each channel, a wide spectrum of colors can be represented. This is the color model typically used by computer screens, digital cameras, and scanners [47].
@@ 7.5.2 HSV and HSI Color Models
While the RGB color model is effective for many tasks, it does not always align with human perception of color. The HSV (Hue-Saturation-Value) and HSI (Hue-Saturation-Intensity) color models are designed to be more perceptually intuitive, allowing for better manipulation in tasks like color-based segmentation and enhancement [58].
@@ 21 V = hsv_image[2, :, :]
  • These models are also used in color pickers in graphic design software [47].
@@ 7.5.3 YCbCr Color Model
The YCbCr color model is commonly used in image and video compression, such as in JPEG and MPEG formats [47]. This model separates the luminance (Y) component from the chrominance components (Cb and Cr), which makes it more efficient for compression, as the human eye is more sensitive to changes in brightness (luminance) than to changes in color (chrominance).
@@
  • Cr (red-difference chrominance) represents the difference between the red channel and the luminance [58].
@@ 20 Cr = ycbcr_image[2, :, :]
  • The YCbCr model is widely used in image and video compression algorithms like JPEG and MPEG [47].
  • It is efficient for compression because the chrominance channels can be subsampled without significantly affecting image quality [47].
@@ 8.1 Spatial Domain Enhancement
Image enhancement in the spatial domain involves directly manipulating the pixel values of an image to improve its visual quality or make specific features more discernible. The spatial domain refers to the image plane itself, and enhancements are achieved by modifying the intensity of pixels or groups of pixels [58]. Several techniques exist for image enhancement in the spatial domain, including histogram equalization [151], contrast stretching [214], and histogram matching [180].
@@ 8.1.1 Histogram Equalization
Histogram equalization works by transforming the intensity values so that the resulting histogram is approximately uniform across the intensity range. The transformation function is based on the cumulative distribution function (CDF) of the image histogram [46].
@@
The following Python code demonstrates how to perform histogram equalization using Pillow and NumPy [36, 199].
@@ 8.1.2 Contrast Stretching
Contrast stretching, also known as dynamic range expansion, is a simple and effective technique used to enhance the contrast of an image by stretching the range of pixel intensity values. Unlike histogram equalization, which redistributes the intensities nonlinearly, contrast stretching is typically a linear operation that maps the input intensity values to cover the full output range (e.g., 0 to 255 for 8-bit images) [214].
@@
Nonlinear contrast stretching applies transformations that are not linear, such as logarithmic or exponential functions. This is useful when the image contains high-contrast areas that should be enhanced or suppressed differently across different intensity ranges [196].
@@ 8.1.3 Histogram Matching
Histogram matching, also known as histogram specification, is a technique that adjusts the intensity distribution of an input image so that its histogram matches that of a target image [180]. Unlike histogram equalization, which enhances contrast globally, histogram matching allows you to match the intensity distribution of one image to that of another, which can be useful in applications like image comparison, object recognition, or aesthetic adjustments [58].
@@ 8.3 Frequency Domain Enhancement
Image enhancement in the frequency domain involves transforming an image from the spatial domain to the frequency domain using the Fourier transform. Once in the frequency domain, various filtering techniques can be applied to manipulate specific frequency components of the image. This can be useful for noise reduction, edge enhancement, and other forms of image improvement [58, 183]. In this section, we will explore the 2D Discrete Fourier Transform (DFT) and how low-pass and high-pass filtering techniques are used in the frequency domain for image enhancement. We will also discuss different filter design strategies.
@@ 8.3.1 Fourier Transform for Image Enhancement
The Fourier transform is a fundamental tool in image processing for converting an image from the spatial domain (where pixel intensities are represented by their position) to the frequency domain (where the image is represented by its frequency components). The 2D Discrete Fourier Transform (DFT) is used to analyze how image intensities change over space [191].
@@
The DFT allows us to analyze an image in terms of its low-frequency components (which represent smooth variations in intensity) and high-frequency components (which represent rapid changes like edges and fine details) [58].
@@ 8.3.2 Low-pass and High-pass Filtering
Filtering in the frequency domain involves modifying the frequency components of an image to achieve a desired enhancement effect. Two common types of filters are low-pass filters and high-pass filters [58].
@@
A low-pass filter allows low-frequency components of the image to pass while attenuating high-frequency components. Low frequencies represent smooth variations in intensity, so low-pass filtering is often used for noise reduction and blurring [121].
@@
A high-pass filter does the opposite of a low-pass filter: it attenuates low-frequency components and allows high-frequency components to pass. High-pass filtering is commonly used for edge detection and sharpening because edges in an image correspond to rapid changes in intensity, which are represented by high-frequency components [121].
@@ 8.3.3 Filter Design in Frequency Domain
Different types of filters can be designed in the frequency domain to meet specific image enhancement goals. Common filters include ideal, Butterworth, and Gaussian filters. Each of these filters has different characteristics in terms of their frequency response and the way they transition between passing and attenuating frequencies [58].
@@
Butterworth filters have a smoother transition between pass and stop bands compared to ideal filters. They do not have a sharp cutoff, which helps reduce ringing artifacts [178]. The transfer function of a Butterworth filter is given by:
@@
Gaussian filters are widely used in image processing due to their smooth frequency response and minimal artifacts [80]. The transfer function of a Gaussian filter is:
@@ 8.4 Image Smoothing and Sharpening
Image processing often involves enhancing or improving the quality of an image by either reducing noise or enhancing important features such as edges [58]. In this section, we will discuss two fundamental techniques in image processing: smoothing and sharpening. Smoothing is primarily used to reduce noise, while sharpening enhances the edges and details in an image.
@@ 8.4.1 Smoothing Filters
There are several types of smoothing filters, each with its specific application and properties. The most common ones include mean, median, and bilateral filters [61, 48].
@@ 8.4.2 Sharpening Filters
Sharpening filters enhance the edges and fine details of an image, making it appear crisper and more detailed. These filters highlight regions in an image where there are abrupt intensity changes, such as edges between objects. This is useful in applications like edge detection, object recognition, and image enhancement [58].
Some of the most commonly used sharpening filters include the Laplacian, Sobel, and Prewitt operators [143, 27].
@@ 1. Laplacian Filter:
Laplacian $\text{Laplacian Kernel}=\begin{bmatrix}0&-1&0\\ -1&4&-1\\ 0&-1&0\end{bmatrix}$
@@ 2. Sobel and Prewitt Operators:
Sobel Horizontal $\text{Sobel Horizontal Kernel}=\begin{bmatrix}-1&0&1\\ -2&0&2\\ -1&0&1\end{bmatrix}$
Sobel Vertical $\text{Sobel Vertical Kernel}=\begin{bmatrix}-1&-2&-1\\ 0&0&0\\ 1&2&1\end{bmatrix}$
@@
Image restoration and reconstruction are key areas of digital image processing, focused on improving the visual quality of images that have been degraded or corrupted. Unlike image enhancement, which emphasizes improving the overall appearance of an image, restoration focuses on reversing known degradation processes to recover the original image. In many cases, restoration involves addressing issues such as noise, blur, and distortions caused by imperfections in the imaging system or during transmission and storage [58].
@@ 9.1 Image Degradation Models
An important part of image restoration is understanding the causes of degradation and developing mathematical models to describe them. Degradation may occur during the acquisition process, when an image is captured using a camera or sensor, during transmission (such as in communication systems), or during storage (such as compression artifacts or data corruption) [146].
@@ 9.1.1 Modeling of Degradation Processes
Image degradation can be represented mathematically by a model that describes how the original image $f(x,y)$ is altered during the acquisition or transmission process [9]. A typical degradation model is represented as:
@@
Blur is one of the most common forms of degradation, often caused by motion, defocus, or atmospheric disturbances [221]. The degradation function $h(x,y)$ typically represents the blurring filter applied to the image. For example:
@@ 9.2 Noise Models
Noise is another common form of image degradation that occurs during image acquisition or transmission. Noise can be introduced by a variety of factors, including sensor limitations, electronic interference, and compression artifacts. Different types of noise have different characteristics and impact the image quality in different ways [19].

### v2

@@ 7.1.2 Image Elements
The building blocks of any digital image are its pixels, which form a grid-like structure representing the image. Each pixel is a tiny square (or sometimes other shapes) that contains a value corresponding to the brightness (in grayscale images) or color (in color images) at that point in the image. The quality and detail of an image depend on two important factors: spatial resolution and grayscale resolution [gonzalez2002digital].
@@
  • In an 8-bit grayscale image, each pixel can have one of 256 possible intensity values (ranging from 0 to 255) [orchard1991color].
@@
  • In a 24-bit color image, each pixel has 8 bits per channel (red, green, and blue), allowing for $256^{3}=16,777,216$ different colors.
The higher the bit depth, the more shades or colors an image can represent, resulting in smoother gradients and more realistic visual representations [heckbert1982color].
@@ 7.2 Python Example: Working with Grayscale Images
Python offers excellent libraries for digital image processing, such as Pillow [clark2015pillow], Matplotlib [hunter2007matplotlib], and OpenCV [bradski2000opencv]. To begin with, let’s work with grayscale images, which are simpler to understand since they only contain intensity information.
@@ 7.4.1 Image Sampling
Image sampling is the process of converting a continuous image into a digital format by taking measurements, or samples, of the image at regular intervals. These samples correspond to pixels in the digital image, and each pixel represents the intensity (or color) of the image at a specific point [gonzalez2002digital].
In the continuous world, images are represented by a function $f(x,y)$, where $x$ and $y$ denote the spatial coordinates, and $f(x,y)$ represents the intensity at that point. When we sample the image, we select a grid of points at intervals and store the intensity values at those locations. This sampled image becomes a discrete grid of pixels, and the spacing between the samples is referred to as the sampling rate or resolution [shannon1949communication].
@@
The sampling rate, or the distance between adjacent samples, has a significant impact on the quality and detail of the digital image. A higher sampling rate (denser grid of pixels) captures more detail from the original image, while a lower sampling rate results in a loss of detail and potential aliasing effects [candes2008introduction].
  • High Sampling Rate: When the sampling rate is high, the image closely resembles the original continuous image. Fine details and sharp edges are preserved [gonzalez2002digital].
@@
The Nyquist Sampling Theorem provides a guideline for selecting an appropriate sampling rate to avoid aliasing [shannon1949communication]. According to this theorem, the sampling rate must be at least twice the highest frequency present in the image to accurately capture all details [oppenheim2010dsp]. This minimum rate is known as the Nyquist rate. If the sampling rate is lower than the Nyquist rate, aliasing occurs, where high-frequency components appear as lower-frequency artifacts in the image [candes2008introduction].
@@ 7.4.2 Image Quantization
After sampling an image in the spatial domain, the next step is to quantize its intensity values. Quantization is the process of mapping the continuous range of intensity values in the original image to a limited set of discrete values in the digital image [gonzalez2002digital]. In practice, each pixel’s intensity in the digital image is represented by a fixed number of bits, which limits the number of possible intensity levels.
@@
  • 8-bit Image: Each pixel can have one of 256 possible intensity levels (from 0 to 255). This is common for grayscale images [kimpe2007increasing].
  • 16-bit Image: Each pixel can have one of 65,536 possible intensity levels (from 0 to 65535), allowing for much finer intensity gradations [guo2009high].
@@
Quantization introduces an error known as quantization error, which is the difference between the actual continuous intensity values and the discrete levels assigned during quantization. This error can manifest as noise or distortion in the image, especially when using a low bit depth [heckbert1982color].
@@ 7.5 Color Spaces
In digital image processing, color spaces are used to represent colors in a format that can be processed and analyzed by computers. Different color spaces are suited for different applications, from basic image display to more complex tasks like image compression, segmentation, and enhancement [gonzalez2002digital]. The choice of color space affects how easily certain operations can be performed on the image, and how colors are interpreted. In this section, we will cover several commonly used color models, including RGB, HSV/HSI, and YCbCr.
@@ 7.5.1 RGB Color Model
The RGB color model is one of the most widely used color models in digital image processing. It represents colors by combining three primary color channels: Red (R), Green (G), and Blue (B). Each pixel in an image has a value for these three channels, and by varying the intensity of each channel, a wide spectrum of colors can be represented. This is the color model typically used by computer screens, digital cameras, and scanners [duan2020video].
@@ 7.5.2 HSV and HSI Color Models
While the RGB color model is effective for many tasks, it does not always align with human perception of color. The HSV (Hue-Saturation-Value) and HSI (Hue-Saturation-Intensity) color models are designed to be more perceptually intuitive, allowing for better manipulation in tasks like color-based segmentation and enhancement [gonzalez2002digital].
@@ 21 V = hsv_image[2, :, :]
  • These models are also used in color pickers in graphic design software [duan2020video].
@@ 7.5.3 YCbCr Color Model
The YCbCr color model is commonly used in image and video compression, such as in JPEG and MPEG formats [duan2020video]. This model separates the luminance (Y) component from the chrominance components (Cb and Cr), which makes it more efficient for compression, as the human eye is more sensitive to changes in brightness (luminance) than to changes in color (chrominance).
@@
  • Cr (red-difference chrominance) represents the difference between the red channel and the luminance [gonzalez2002digital].
@@ 20 Cr = ycbcr_image[2, :, :]
  • The YCbCr model is widely used in image and video compression algorithms like JPEG and MPEG [duan2020video].
  • It is efficient for compression because the chrominance channels can be subsampled without significantly affecting image quality [duan2020video].
@@ 8.1 Spatial Domain Enhancement
Image enhancement in the spatial domain involves directly manipulating the pixel values of an image to improve its visual quality or make specific features more discernible. The spatial domain refers to the image plane itself, and enhancements are achieved by modifying the intensity of pixels or groups of pixels [gonzalez2002digital]. Several techniques exist for image enhancement in the spatial domain, including histogram equalization [pizer1987adaptive], contrast stretching [yang2006image], and histogram matching [shen2007image].
@@ 8.1.1 Histogram Equalization
Histogram equalization works by transforming the intensity values so that the resulting histogram is approximately uniform across the intensity range. The transformation function is based on the cumulative distribution function (CDF) of the image histogram [drew2000computing].
@@
The following Python code demonstrates how to perform histogram equalization using Pillow and NumPy [clark2015pillow, van2011numpy].
@@ 8.1.2 Contrast Stretching
Contrast stretching, also known as dynamic range expansion, is a simple and effective technique used to enhance the contrast of an image by stretching the range of pixel intensity values. Unlike histogram equalization, which redistributes the intensities nonlinearly, contrast stretching is typically a linear operation that maps the input intensity values to cover the full output range (e.g., 0 to 255 for 8-bit images) [yang2006image].
@@
Nonlinear contrast stretching applies transformations that are not linear, such as logarithmic or exponential functions. This is useful when the image contains high-contrast areas that should be enhanced or suppressed differently across different intensity ranges [trongtirakul2019non].
@@ 8.1.3 Histogram Matching
Histogram matching, also known as histogram specification, is a technique that adjusts the intensity distribution of an input image so that its histogram matches that of a target image [shen2007image]. Unlike histogram equalization, which enhances contrast globally, histogram matching allows you to match the intensity distribution of one image to that of another, which can be useful in applications like image comparison, object recognition, or aesthetic adjustments [gonzalez2002digital].
@@ 8.3 Frequency Domain Enhancement
Image enhancement in the frequency domain involves transforming an image from the spatial domain to the frequency domain using the Fourier transform. Once in the frequency domain, various filtering techniques can be applied to manipulate specific frequency components of the image. This can be useful for noise reduction, edge enhancement, and other forms of image improvement [gonzalez2002digital, singh2014various]. In this section, we will explore the 2D Discrete Fourier Transform (DFT) and how low-pass and high-pass filtering techniques are used in the frequency domain for image enhancement. We will also discuss different filter design strategies.
@@ 8.3.1 Fourier Transform for Image Enhancement
The Fourier transform is a fundamental tool in image processing for converting an image from the spatial domain (where pixel intensities are represented by their position) to the frequency domain (where the image is represented by its frequency components). The 2D Discrete Fourier Transform (DFT) is used to analyze how image intensities change over space [sundararajan2001discrete].
@@
The DFT allows us to analyze an image in terms of its low-frequency components (which represent smooth variations in intensity) and high-frequency components (which represent rapid changes like edges and fine details) [gonzalez2002digital].
@@ 8.3.2 Low-pass and High-pass Filtering
Filtering in the frequency domain involves modifying the frequency components of an image to achieve a desired enhancement effect. Two common types of filters are low-pass filters and high-pass filters [gonzalez2002digital].
@@
A low-pass filter allows low-frequency components of the image to pass while attenuating high-frequency components. Low frequencies represent smooth variations in intensity, so low-pass filtering is often used for noise reduction and blurring [makandar2015image].
@@
A high-pass filter does the opposite of a low-pass filter: it attenuates low-frequency components and allows high-frequency components to pass. High-pass filtering is commonly used for edge detection and sharpening because edges in an image correspond to rapid changes in intensity, which are represented by high-frequency components [makandar2015image].
@@ 8.3.3 Filter Design in Frequency Domain
Different types of filters can be designed in the frequency domain to meet specific image enhancement goals. Common filters include ideal, Butterworth, and Gaussian filters. Each of these filters has different characteristics in terms of their frequency response and the way they transition between passing and attenuating frequencies [gonzalez2002digital].
@@
Butterworth filters have a smoother transition between pass and stop bands compared to ideal filters. They do not have a sharp cutoff, which helps reduce ringing artifacts [selesnick1998generalized]. The transfer function of a Butterworth filter is given by:
@@
Gaussian filters are widely used in image processing due to their smooth frequency response and minimal artifacts [ito2000gaussian]. The transfer function of a Gaussian filter is:
@@ 8.4 Image Smoothing and Sharpening
Image processing often involves enhancing or improving the quality of an image by either reducing noise or enhancing important features such as edges [gonzalez2002digital]. In this section, we will discuss two fundamental techniques in image processing: smoothing and sharpening. Smoothing is primarily used to reduce noise, while sharpening enhances the edges and details in an image.
@@ 8.4.1 Smoothing Filters
There are several types of smoothing filters, each with its specific application and properties. The most common ones include mean, median, and bilateral filters [griffin2000mean, elad2002origin].
@@ 8.4.2 Sharpening Filters
Sharpening filters enhance the edges and fine details of an image, making it appear crisper and more detailed. These filters highlight regions in an image where there are abrupt intensity changes, such as edges between objects. This is useful in applications like edge detection, object recognition, and image enhancement [gonzalez2002digital].
Some of the most commonly used sharpening filters include the Laplacian, Sobel, and Prewitt operators [paris2011local, chaple2015comparisions].
@@ 1. Laplacian Filter:
$\text{Laplacian Kernel}=\begin{bmatrix}0&-1&0\\ -1&4&-1\\ 0&-1&0\end{bmatrix}$
@@ 2. Sobel and Prewitt Operators:
$\text{Sobel Horizontal Kernel}=\begin{bmatrix}-1&0&1\\ -2&0&2\\ -1&0&1\end{bmatrix}$
$\text{Sobel Vertical Kernel}=\begin{bmatrix}-1&-2&-1\\ 0&0&0\\ 1&2&1\end{bmatrix}$
@@
Image restoration and reconstruction are key areas of digital image processing, focused on improving the visual quality of images that have been degraded or corrupted. Unlike image enhancement, which emphasizes improving the overall appearance of an image, restoration focuses on reversing known degradation processes to recover the original image. In many cases, restoration involves addressing issues such as noise, blur, and distortions caused by imperfections in the imaging system or during transmission and storage [gonzalez2002digital].
@@ 9.1 Image Degradation Models
An important part of image restoration is understanding the causes of degradation and developing mathematical models to describe them. Degradation may occur during the acquisition process, when an image is captured using a camera or sensor, during transmission (such as in communication systems), or during storage (such as compression artifacts or data corruption) [pei2019effects].
@@ 9.1.1 Modeling of Degradation Processes
Image degradation can be represented mathematically by a model that describes how the original image $f(x,y)$ is altered during the acquisition or transmission process [baird2007state]. A typical degradation model is represented as:
@@
Blur is one of the most common forms of degradation, often caused by motion, defocus, or atmospheric disturbances [zhou2017classification]. The degradation function $h(x,y)$ typically represents the blurring filter applied to the image. For example:
@@ 9.2 Noise Models
Noise is another common form of image degradation that occurs during image acquisition or transmission. Noise can be introduced by a variety of factors, including sensor limitations, electronic interference, and compression artifacts. Different types of noise have different characteristics and impact the image quality in different ways [boyat2015review].